An automobile, moving to fast on a horizontal stretch of mountain road, slides off the roads, falling into deep snow 43.9 m below the road and 87.7 m beyond the edge of the road
A) how long did the auto take to fall?
1
Expert's answer
2020-07-06T15:14:49-0400

The kinematic law of motion under the gravity force is given be the following equation:


"d = v_0t + \dfrac{gt^2}{2}"

where "d = 43.9 m" is traveled distance, "v_0 = 0m\/s" is vertical initial speed, "g = 9.8 m\/s^2" and "t" is a time of falling.

Expressing "t" and substituting numerical values, get:


"t = \sqrt{\dfrac{2d}{g}} = \sqrt{\dfrac{2\cdot 43.9}{9.8}} \approx 3.17s"

Answer. 3.17 s